OpenStudy (anonymous):

how many significant figures are in the measurement 40,500 mg

OpenStudy (anonymous):

3. Here's how it works: if we're talking about numbers without decimal places, the last number from left to right that's not zero is the last significant digit. Let's count here: 4 (1 digit) 0 (2 digits) 5 (3 digits), and the rest of the zeros don't count

OpenStudy (cuanchi):

The expression as 40,500 mg is confusing, @Rodriguez47 is right, but some one else can come with other argument saying that that zeros are significant. To avoid confusion some authors use a decimal place at the end of the number as this 40,500. to indicate that the last to zeros are significant. To avoid confusion the most practical way to express the number with the correct number of significant figures will be in the scientific notation. 4.0500 x 10^4 = 5 significant figures 4.050 x 10^4 = 4 significant figures 4.05 x 10^4 = 3 significant figures